# Think, Reflect, Create: Metacognitive Learning for Zero-Shot Robotic Planning with LLMs

### v1

@@ Abstract
While large language models (LLMs) have shown great potential across various domains, their applications in robotics remain largely limited to static, prompt-based behaviors and still face challenges in handling complex tasks under zero-shot or few-shot settings. Inspired by human metacognitive learning and creative problem-solving, we address this limitation by exploring a fundamental research question: Can LLMs be empowered with metacognitive capabilities to reason, reflect, and create, thereby enhancing their ability to perform robotic tasks with minimal demonstrations? In this paper, we present an early-stage framework that integrates metacognitive learning into LLM-powered multi-robot collaboration. The proposed framework equips the LLM-powered robotic agents with a skill decomposition and self-reflection mechanism that identifies modular skills from prior tasks, reflects on failures in unseen task scenarios, and synthesizes effective new solutions. Experimental results show that our metacognitive-learning-empowered LLM framework significantly outperforms existing baselines. Moreover, we observe that the framework is capable of generating solutions that differ from the ground truth yet still successfully complete the tasks. These exciting findings support our hypothesis that metacognitive learning can foster creativity in robotic planning.
Think, Reflect, Create: Metacognitive Learning for Zero-Shot Robotic Planning with LLMs
Wenjie Lin and Jin Wei-Kocsis††thanks: Corresponding author. Department of Computer and Information Technology Purdue University {lin1790, kocsis0}@purdue.edu
1 Introduction
@@
In recent years, large language models (LLMs) have emerged as powerful reasoning engines capable of performing complex planning, decision-making, and knowledge-intensive tasks across various application domains. These successes have sparked growing interest in applying LLMs to robotic operations, enabling robots to understand instructions, generate executable action sequences, and generalize across diverse and novel scenarios (Wang et al., 2024; Liu et al., 2024; Jin et al., 2024; Tan et al., 2024; Cheng et al., 2024; Chen et al., 2025). Recent research on LLM-powered robotic operations can be broadly categorized into three main directions: 1) generating robot plans by prompting LLMs with task instructions (Mandi et al., 2024), 2) enabling embodied reasoning by integrating LLMs with multimodal perception systems (Aissi et al., 2025), and 3) synthesizing robot control codes from natural language commands (Liang et al., 2023). While these approaches demonstrate the potential of LLMs to advance robotic operations, most remain limited to static prompting, which restricts their performance in complex tasks in zero-shot or few-shot settings.
@@
Inspired by human metacognitive learning and its impact on enabling creative skills (Hargrove, 2013; Schuster et al., 2020), we address this limitation by exploring a fundamental question: Can LLMs be endowed with metacognitive capabilities to reason, reflect, and create, thereby enhancing their ability to perform robotic tasks with minimal demonstrations? In this paper, we present an early-stage framework that integrates metacognitive learning into LLM-powered multi-robot collaboration. The proposed system equips the LLM with a skill decomposition and self-reflection mechanism that identifies modular skills from prior tasks, reflects on failures in novel scenarios, and synthesizes effective new solutions. Our work in this paper has three key contributions: (1) To the best of our knowledge, this is the first work to explore integrating metacognitive learning into LLM-equipped robot operations, to support both reliable performance and creative problem-solving; (2) We propose a metacognitive learning framework that enables the LLM-powered robotic agents decompose modular skills, reflect on task failures, and synthesize effective new solutions; and (3) We validate our framework on the RoCo benchmark, where it significantly outperforms baselines and sometimes generates successful solutions that deviate from the ground truth, which supports the hypothesis that metacognitive learning can foster reliable and creative robotic planning.
2 Problem Formulation
| Move Rope | Arrange Cabinet | Make Sandwich | ||
@@ 2 Problem Formulation
| Our framework | Task Success Rate | 0.76 0.10 | 0.95 0.05 | 0.95 0.05 |
| Environment Steps, Replan Attempts | 2.0, 2.4 | 4.0, 1.7 | 9.4, 1.8 |
We consider a cooperative multi-agent manipulation setting, where multiple LLM-powered robot agents collaborate to complete tasks over a finite time horizon. Each agent operates within its own observation space and needs to coordinate with other agents to achieve a shared task objective. At each time step, each agent receives a prompt and outputs , where denotes the resulting arm motion plan, represents the agent-specific task description that includes goals and constraints, is its current observation, and is the metacognition-informed input.
The generation of effective is guided by our proposed metacognitive learning module, which empowers the LLM to decompose prior task completions into modular skills, synthesize arm motion plans for unseen task scenarios, reflect on planning generation failures, and iteratively produce effective and potentially creative solutions. In our current implementation, we adopted the validation mechanism in (Mandi et al., 2024) to detect planning generation failures and trigger the self-reflection process.
3 Methodology
The proposed metacognitive learning module is illustrated in Fig. 1. As shown in the figure, the module comprises three key components: (1) modular skill set construction, (2) metacognitive inference, and (3) self-reflection.
In the modular skill set construction component, the LLM is guided by the metacognition-informed input to construct a library of exemplars from previously completed tasks. Each exemplar maps identified robot manipulation skills to a successful task execution and includes a representative scene along with a one-shot demonstration of an effective action plan. The LLM then extracts fine-grained modular skills from each exemplar, clusters similar skills to reduce redundancy, and organizes them into a reusable library of transferable modular robot manipulation skills and their associated exemplars.
In the metacognitive inference component, based on the library of modular robot manipulation skills and associated exemplars, the task description , and the current observation , the LLM is guided by to reason about the necessary modular skills for the new task. Using the identified skills and their associated exemplars, the LLM synthesizes arm motion plans for the robot agent .
The self-reflection component is activated when the arm motion plans synthesized during metacognitive inference do not pass the validation check. It guides the LLM to reflect on failure feedback to identify modular skills that are missing or need refinement. Based on these insights, the LLM retrieves the corresponding exemplars and synthesizes revised arm motion plans.
4 Experiments
We validate the performance of the proposed metacognitive learning module in enabling LLM-enabled robotic agents to complete complex multi-robot collaboration tasks under zero-shot settings. In this initial stage of work, we conduct experiments using the RoCo benchmark (Mandi et al., 2024), focusing on three challenging tasks: Move Rope, Arrange Cabinet, and Make Sandwich. In each task, LLM-powered robot agents coordinate via structured textual prompts that include metacognition-informed inputs and execute actions in environments with obstacles and spatial constraints. An embedded validation mechanism detects planning failures, such as those caused by collisions or inverse kinematics (IK) infeasibility, and triggers the self-reflection process. All experiments are conducted on a machine with an NVIDIA A100 GPU to support efficient inference using the LLaMA 3.1-70B model. We note that LLaMA is adopted instead of GPT-4 in our framework to align with our long-term objective of developing an open-source LLM framework that can be widely adopted by the robotics and AI research communities.
@@ 4 Experiments
We compare our framework against two baselines. (1) Central Plan: an oracle LLM-based planner with access to the full environment state, task description, and capabilities of all robots. It generates a joint centralized plan without accounting for information asymmetry. (2) RoCo+GPT-4: the state-of-the-art multi-robot collaboration framework proposed in (Mandi et al., 2024), which uses GPT-4 but does not incorporate metacognition-informed input. To ensure a fair performance comparison, we follow RoCoBench (Mandi et al., 2024) and use the same evaluation metrics: (1) Task Success Rate, which measures the percentage of successful task completions within a fixed number of rounds (we also use over 20 rounds in our experiements); (2) Environment Steps, defined as the average number of steps taken in successful runs; and (3) Replan Attempts, which refers to the average number of replan attempts across all runs.
Experiment Findings
@@
Reliability Enhancement: The evaluation results comparing our proposed framework with the two baselines across the three tasks are shown in Table 1. In the Move Rope task, which is the most challenging task in the experiments, our framework achieves a success rate of , representing a improvement over RoCo+GPT-4 and a gain over Central Plan. It also reduces environment steps in successful runs to , compared to for RoCo+GPT-4 and for Central Plan, and lowers replan attempts to , compared to for RoCo+GPT-4 and for Central Plan. These results indicate more effective coordination and faster convergence under challenging spatial constraints. For the Arrange Cabinet task, our framework achieves a success rate of , outperforming RoCo by and the Central Plan by . Additionally, it requires the same number of environment steps as the Central Plan while requiring fewer replan attempts than both baselines, suggesting improved planning robustness and reduced reliance on corrective execution. For the Make Sandwich task, which involves long-horizon planning and strict stacking constraints, our framework achieves performance comparable to Central Plan. In comparison to the RoCo+GPT-4 baseline, it achieves a higher success rate and reduces the required environment steps, highlighting its ability to generalize to structurally complex tasks with minimal planning overhead, despite a slight increase in replan attempts. These results validate the effectiveness of our framework in advancing the capabilities of LLM-powered robot agents for completing complex tasks under zero-shot settings. They suggest that the proposed metacognitive learning module enables LLM-equipped robot agents to adaptively and proactively reason about and reflect on spatial, temporal, and structural challenges.
@@
Creativity Cultivation: In the experiments, we observe that our framework is capable of generating solutions that differ from the ground truth yet still successfully complete the tasks. These exciting findings support our hypothesis that metacognitive learning can cultivate creativity in robotic planning. Due to space limitations, we describe one representative case. In the Move Rope task, two robot agents collaboratively grasp the ends of a rope, maneuver it over an obstacle wall, and place it into a designated groove. In the ground-truth plan, the robots grasp the two ends of the rope to complete the task. In contrast, after initial plan generation failures due to collision or IK infeasibility, our framework generates an alternative motion plan in which one robot grasps the rope slightly inward instead of at the end. Although this alternative plan deviates from the ground truth, it effectively shortens the trajectory, improves the spatial separation between the robot arms, and significantly reduces the risk of collision and IK errors.
| Move Rope | Arrange Cabinet | Make Sandwich | |
| Reflection Success Rate | 0.44 | 1.00 | 0.70 |
@@
Metacognitive Self-Reflection Analysis: As shown in Fig. 1, the self-reflection component is one of the three key elements of our proposed metacognitive learning module. To evaluate its practical impact, we introduce an additional metric in our experiments, the reflection success rate, defined as the proportion of successful plan regenerations among all reflection attempts within task rounds that ultimately succeeded. The results, summarized in Table 2. As shown in the table, in the Arrange Cabinet task, the framework achieves a perfect reflection success rate of , indicating that every initial failure was successfully recovered through self-reflection. In the Make Sandwich task, which involves long-horizon dependencies and stacking constraints, the framework recovers from of failed plans via self-reflection. Even in the most challenging Move Rope task, characterized by tight spatial coordination, the framework achieves a reflection success rate, underscoring the role of self-reflection in enabling meaningful plan recovery under physical constraints. These findings highlight the critical role of the self-reflection component within our metacognitive learning module. They also show that the proposed framework not only supports LLM-powered agents in initial plan generation but also enables them to proactively reason about and reflect on execution failures, providing the essential capability of reliable recovery and adaptation in zero-shot settings.
5 Conclusions

### v2

@@ Abstract
While large language models (LLMs) have shown great potential across various domains, their applications in robotics remain largely limited to static prompt-based behaviors and still face challenges in complex tasks under zero-shot or few-shot settings. Inspired by human metacognitive learning and creative problem-solving, we address this limitation by exploring a fundamental question: Can LLMs be empowered with metacognitive capabilities to reason, reflect, and create, thereby enhancing their ability to perform robotic tasks with minimal demonstrations? In this paper, we present a framework that integrates metacognitive learning into LLM-powered multi-robot collaboration. The system equips the LLM-powered robotic agents with a skill decomposition and self-reflection mechanism that identifies modular skills from prior tasks, reflects on failures in unseen task scenarios, and synthesizes effective new solutions. We propose a more challenging robotic benchmark task and evaluate our framework on the existing benchmark and the novel task. Experimental results show that our metacognitive learning framework significantly outperforms existing baselines. Moreover, we observe that the framework can generate solutions that differ from the ground truth yet still successfully complete the tasks. These findings support our hypothesis that metacognitive learning can foster creativity in robotic planning.
Introduction
@@
In recent years, large language models have emerged as powerful reasoning engines capable of performing complex planning, decision-making, and knowledge-intensive tasks across various application domains. These successes have sparked growing interest in applying LLMs to robotic operations, enabling robots to understand instructions, generate executable action sequences, and generalize across diverse and novel scenarios (Wang et al. 2024; Liu et al. 2024; Jin et al. 2024; Tan et al. 2024; Cheng et al. 2024; Chen et al. 2025). Recent research on LLM-powered robotic operations can be broadly categorized into three main directions: 1) generating robot plans by prompting LLMs with task instructions (Mandi, Jain, and Song 2024), 2) enabling embodied reasoning by integrating LLMs with multimodal perception systems (Aissi et al. 2025), and 3) synthesizing robot control codes from natural language commands (Liang et al. 2023). While these approaches demonstrate the potential of LLMs to advance robotic operations, most existing approaches lack mechanisms for metacognitive reasoning or dynamic adaptation after failures, restricting their effectiveness in complex zero-shot or few-shot tasks.
@@
Inspired by human metacognitive learning and its impact on enabling creative skills (Hargrove 2013; Schuster et al. 2020), we address this limitation by exploring a fundamental question: Can LLMs be endowed with metacognitive capabilities to reason, reflect, and create, thereby enhancing their ability to perform robotic tasks with minimal demonstrations? In this paper, we present an early-stage framework that integrates metacognitive learning into LLM-powered multi-robot collaboration. The proposed system equips the LLM with skill decomposition, metacognitive inference, and self-reflection mechanisms that identify modular skills from prior tasks, reflect on failures in novel scenarios, and synthesize effective new solutions. Our work makes three key contributions: (1) To the best of our knowledge, this is the first work to explore integrating metacognitive learning into LLM-equipped robot operations, to support both reliable performance and creative problem-solving; (2) We propose a metacognitive learning framework that enables the LLM-powered robotic agents decompose modular skills, metacognitive inference, reflect on task failures, and synthesize effective new solutions; and (3) We develop a novel robotic benchmark task and validate our framework on both our new and existing benchmark tasks, where it significantly outperforms baselines and sometimes generates valid solutions that deviate from the ground truth, which supports the hypothesis that metacognitive learning can foster reliable and creative robotic planning.
The rest of the paper is organized as follows. We first present related work and problem formulation, and describe our metacognitive learning framework. Experiments and results follow, along with discussions and limitations. We conclude in the final section.
Related Work
LLM-powered Robotic Systems:
LLMs have demonstrated great potential in robotic planning and reasoning. Recent advances include prompt-based planning (Mandi, Jain, and Song 2024), multimodal perception integration (Tan et al. 2024; Aissi et al. 2025), and direct synthesis of robot control code from natural language (Liang et al. 2023; Chen et al. 2025). Specifically, RoCo (Mandi, Jain, and Song 2024) proposed a dialectic approach for LLM-powered multi-robot collaboration. VIPER (Aissi et al. 2025) was introduced by integrating LLMs with visual inputs for sequential decision-making, but without reflection or creative capabilities. Researchers addressed long-horizon planning through multimodal skill learning (Tan et al. 2024) and multi-view world model (Chen et al. 2025), respectively, yet neither incorporated metacognitive reasoning. Direct policy generation was enabled from natural language (Liang et al. 2023) but relied on predefined tasks without adaptive reflection.
Metacognitive Learning:
Metacognitive learning, focused on regulation of cognitive processes, has proven beneficial to enhance creativity and problem-solving skills in humans (Hargrove 2013; Schuster et al. 2020), but its application in robotics remains limited. Prior studies highlight its benefits for learning efficiency, adaptability, and creative skill development in education contexts (Schuster et al. 2020; Hargrove 2013), providing a foundation for our work. Although some efforts in LLMs and robotics explore self-reflection using reinforcement or iterative learning, they often lack explicit integration of metacognitive reasoning. Our work addresses this gap by explicitly embedding metacognitive components into LLM-driven robotic collaboration.
Creativity in Robotic Planning:
Creativity in robotic task planning often implies the ability to generate effective solutions beyond explicit training data or demonstrations. Prior research on creativity has largely focused on writing or art, emphasizing adaptive generalization or task reformation (Gómez-Rodríguez and Williams 2023). To the best of our knowledge, no existing work focuses on creativity in robotics, where solutions must operate under complex physical and operational constraints. In this work, we take a step toward exploring the capabilities of a metacognition learning-based framework to enable structured creativity in robotic planning. While our work is still in an early stage, it significantly expands the creative potential of robotic systems in zero-shot and minimal-demonstration contexts.
Problem Formulation
In our work, we consider a cooperative multi-agent manipulation setting, where multiple LLM-powered robot agents collaborate to complete tasks over a finite time horizon. Each agent operates within its own observation space and needs to coordinate with other agents to achieve a shared task objective. At each time step, each agent receives a prompt and outputs , where denotes the resulting arm motion plan, represents the agent-specific task description that includes goals and constraints, is its current observation, and is the metacognition-informed input.
The metacognition-informed input is a dynamic guidance signal embedded in the structured prompt provided to each LLM-powered robot agent. It evolves across the three components of the metacognitive learning module and provides essential context that enables the LLM to (1) decompose prior task completions into modular skills, (2) synthesize arm motion plans for unseen task scenarios, (3) reflect on planning generation failures, and (4) iteratively produce effective and potentially creative solutions. More details on the role and structure of at each metacognitive-learning stage will be introduced in the following section.
Methodology
The proposed metacognitive learning framework is illustrated in Figure 1. It comprises three key components: (1) modular skill set construction, (2) metacognitive inference, and (3) self-reflection.
In the modular skill set construction component, the LLM is guided by the metacognition-informed input , which includes retrieved exemplars from previously successful tasks and structured prompt content designed to guide the LLM to support skill abstraction and cluster modular skills. Each exemplar contains a brief task summary, a representative scene, and a one-shot demonstration of an effective action plan. Guided by , the LLM then extracts robot manipulation skills from each exemplar, clusters similar skills that correspond to similar goals to reduce redundancy, and organizes them into a reusable library of transferable modular robot manipulation skills and their associated exemplars.
In the metacognitive inference component, the LLM receives the task description , the current observation , and access to the previously constructed library of modular skills and associated exemplars. The metacognition-informed input serves as a reasoning-oriented guidance signal that provides attention cues to focus the LLM’s inference. Rather than specifying relevant skills directly, helps the LLM reason about which modular skills from the library are applicable to the current task. Using the identified skills and their associated exemplars, the LLM synthesizes arm motion plans for the robot agent .
The self-reflection component is activated when the arm motion plans synthesized during metacognitive inference fail to pass the validation check (e.g. due to collision or inverse kinematics (IK) infeasibility). In our current implementation, we adopted the validation mechanism in (Mandi, Jain, and Song 2024) to detect planning generation failures and trigger the self-reflection process. In this stage, the metacognition-informed input is updated to encode structured failure feedback, including the nature and location of the failure. This updated serves as a reflective prompt that guides the LLM to reason about which modular skills may be missing or misapplied. Guided by these insights, the LLM retrieves relevant exemplars from the skill library and synthesizes a revised arm motion plan to correct the previous failure.
Experiments
We evaluate the performance of the proposed metacognitive learning module in enabling LLM-enabled robotic agents to complete complex multi-robot collaboration tasks under zero-shot settings. We conduct experiments using the RoCoBench (Mandi, Jain, and Song 2024), focusing on four challenging tasks: Move Rope, Arrange Cabinet, Make Sandwich, as well as our newly introduced Install Drywall. Our Install Drywall task represents a significant advancement in multi-robot coordination complexity, requiring continuous safety monitoring, precise spatial alignment, and coordinated load-bearing throughout the task execution process. In all of these tasks, LLM-powered robot agents coordinate via structured textual prompts that include metacognition-informed inputs and execute actions in environments with obstacles and operational constraints. An embedded validation mechanism detects planning failures, such as those caused by collisions or IK infeasibility, and triggers the self-reflection process. All experiments are conducted on a machine with an NVIDIA A100 GPU to support efficient inference using the LLaMA-3.1-70B model and the GPT-4 model for fair comparison.
| Metrics | Move Rope | Arrange Cabinet | Make Sandwich | |
@@
| Ours+LLaMA-3.1 | Task Success Rate | 0.76 0.10 | 0.95 0.05 | 0.95 0.05 |
| Environment Steps, Replan Attempts | 2.0, 2.4 | 4.0, 1.7 | 9.4, 1.8 | |
| Ours+GPT-4 | Task Success Rate | 0.86 0.08 | 1.00 0.00 | 0.95 0.05 |
| Environment Steps, Replan Attempts | 2.1, 1.2 | 4.0, 0.0 | 9.1, 0.0 |
Benchmark
The visualization of each task in this paper is shown in Figure 2.
Install Drywall: A New Multi-Robot Construction Task
To evaluate the metacognitive capabilities of LLM-powered robotic systems in complex scenarios resembling real-world jobsites, we design a novel task for multi-robot collaboration, Install Drywall. In this task, the robot agents must coordinate to pick up and hold the left and right panel handles, lift the panel to a specified height, detect and correct misalignment between the panel and studs on the wall by applying a twist operation, and finally position the panel against the wall with precise parallel alignment. This task elevates the complexity of multi-robot coordination by explicitly requiring advanced and synchronized spatial reasoning and coordinated physical manipulation. Specifically, two robotic agents must collaborate to install a drywall panel onto a wall frame. Specifically, this task incorporates five distinct modular skills: (1) coordinated dual-agent execution, (2) path and collision planning, (3) spatial reasoning and alignment, (4) conditional or context-aware execution, and (5) object manipulation and transfer. Additionally, the Install Drywall task also mandates continuous real-time safety checks, stability monitoring, and precise alignment corrections. Robots must dynamically respond to environmental feedback, reflecting on any encountered failures, and iteratively refining their plans. The task explicitly evaluates the ability of LLM-powered robots to adaptively, and even creatively, generate feasible solutions under challenging zero-shot conditions, thereby effectively validating the robustness and adaptability of our proposed metacognitive learning framework.
RoCoBench
The RoCo benchmark (Mandi, Jain, and Song 2024) is a comprehensive evaluation platform designed explicitly for multi-robot collaboration tasks, addressing a broad spectrum of coordination scenarios. We select the three most challenging and complex tasks from it: Move Rope, Arrange Cabinet, and Make Sandwich. Each task is structured to test specific coordination behaviors, encompassing sequential and parallel task decomposition, varying levels of workspace overlap, and both symmetric and asymmetric observation spaces. The tasks demand sophisticated interaction among robots, including high-level task reasoning, effective information exchange, adaptive sub-task planning, and precise low-level motion planning facilitated by a centralized multi-arm trajectory planner. In our experiment, we leverage the multi-robot collaboration task environment and validation mechanism provided by the RoCoBench, while building our own metacognitive learning-empowered decision-making framework.
@@
We compare our framework against two baselines. (1) Central Plan: an oracle LLM-based planner with access to the full environment state, task description, and capabilities of all robots. It generates a joint centralized plan without accounting for information asymmetry. (2) RoCo+GPT-4: the state-of-the-art multi-robot collaboration framework proposed in (Mandi, Jain, and Song 2024), which uses GPT-4 but does not incorporate metacognition-informed input. To ensure a fair performance comparison, we follow RoCoBench (Mandi, Jain, and Song 2024) and use the same evaluation metrics: (1) Task Success Rate, which measures the percentage of successful task completions within a fixed number of rounds; (2) Environment Steps, defined as the average number of steps taken in successful runs; and (3) Replan Attempts, which refers to the average number of replan attempts across all runs. For fair comparison with RoCo, we use over 20 rounds, and up to 10 steps and 5 replans per run.
| Metrics | Install Drywall | |
| RoCo+GPT-4 | success | 0.62 0.11 |
| step, replan | 5.3, 5.8 | |
| Ours+LLaMA-3.1 | success | 0.95 0.05 |
| step, replan | 4.2, 1.9 | |
| Ours+GPT-4 | success | 1.00 0.00 |
| step, replan | 4.1, 0.0 |
Results
Modular Robotic Skill Library Construction
| Task | Skills | Necessary Modular Skills |
| Install Drywall | synchronized_dual_grasping synchronized_lifting pose_alignment_checking panel_twisting_and_adjustment precise_object_positioning | coordinated_dual_agent_execution spatial_reasoning_and_alignment object_manipulation_and_transfer conditional_or_contextual_execution spaced_path_and_collision_planning |
| Move Rope | synchronized_bimanual_lifting obstacle_avoidance grasp_distance_maintenance | coordinated_dual_agent_execution spatial_reasoning_and_alignment spaced_path_and_collision_planning object_manipulation_and_transfer |
| Arrange Cabinet | door_opening_mechanics multi_arm_synchronization gripper_state_management placement_accuracy | coordinated_dual_agent_execution spaced_path_and_collision_planning object_manipulation_and_transfer conditional_or_contextual_execution |
| Make Sandwich | sequential_object_stacking task_allocation_under_constraints unilateral_reach_limitation_handling action_conflict_avoidance | task_decomposition_and_ordering spatial_reasoning_and_alignment conditional_or_contextual_execution object_manipulation_and_transfer |
To support generalization across diverse manipulation scenarios, we construct a modular robotic skill library by decomposing complex tasks into reusable modular skills. Our library is grounded in the principle of functional modularity, where each modular skill encapsulates linked exemplar subtask plans and can be recombined across tasks. This enables the LLM-based planner to generalize to novel task scenes and adaptively reconfigure plans in a zero-shot setting.
Table 3 summarizes the mapping from high-level tasks to their corresponding low-level skills and the necessary modular skills, which are achieved by modular skill set construction and metacognitive inference components of our proposed module.
Reliability Enhancement
@@ Results
The evaluation results comparing our proposed framework with the two baselines across the three tasks in RoCo are shown in Table 1. In the Move Rope task, which is the most challenging task in the experiments, our framework achieves a success rate of , representing a improvement over RoCo+GPT-4 and a gain over Central Plan. It also reduces environment steps in successful runs to , compared to for RoCo+GPT-4 and for Central Plan, and lowers replan attempts to , compared to for RoCo+GPT-4 and for Central Plan. These results indicate more effective coordination and faster convergence under challenging spatial constraints. For the Arrange Cabinet task, our framework achieves a success rate of , outperforming RoCo by and the Central Plan by . Additionally, it requires the same number of environment steps as the Central Plan while requiring fewer replan attempts than both baselines, suggesting improved planning robustness and reduced reliance on corrective execution. For the Make Sandwich task, which involves long-horizon planning and strict stacking constraints, our framework achieves performance comparable to Central Plan. In comparison to the RoCo+GPT-4 baseline, it achieves a higher success rate and reduces the required environment steps, highlighting its ability to generalize to structurally complex tasks with minimal planning overhead, despite a slight increase in replan attempts. These results validate the effectiveness of our framework in advancing the capabilities of LLM-powered robot agents for completing complex tasks under zero-shot settings. They suggest that the proposed metacognitive learning module enables LLM-equipped robot agents to adaptively and proactively reason about and reflect on spatial, temporal, and structural challenges.
In the Install Drywall task, our framework also demonstrates substantial improvements in both task success rate and efficiency compared to the RoCo+GPT-4 baseline, which is shown in Table 2. Specifically, our framework utilizing LLaMA-3.1 achieves a task success rate of 0.95, representing a significant 53% improvement over RoCo’s success rate of 0.62. Additionally, it reduces both environment steps and replan attempts from 5.3 and 5.8 (baseline) to 4.2 and 1.9, respectively. When employing GPT-4, our framework achieves a high success rate of 1.00 with further reductions in environment steps, which is 4.1, and zero replan attempts. These results validate our framework’s advanced effectiveness in handling complex coordination, real-time adaptability, and precise execution, thus significantly enhancing the reliability and efficiency of robotic task execution under challenging zero-shot conditions.
Creativity Cultivation
@@
In the experiments, we observe that our framework is capable of generating solutions that differ from the ground truth yet still successfully complete the tasks, which is demonstrated in Figure 3. These findings support our hypothesis that metacognitive learning can cultivate structured creativity in robotic planning. To further illustrate this, we provide one representative case. In the Move Rope task, two robot agents need to collaboratively grasp the ends of a rope, maneuver it over an obstacle wall, and place it into a designated groove. In the ground-truth plan, the robots grasp the two ends of the rope to complete the task. In contrast, after initial plan generation failures due to collision or IK infeasibility, our framework generates an alternative motion plan in which one robot grasps the rope slightly inward instead of at the end. Although this alternative plan deviates from the ground truth, it effectively shortens the trajectory, improves the spatial separation between the robot arms, and significantly reduces the risk of collisions and IK errors.
| Reflection Success Rate | LLaMA-3.1 | GPT-4 |
| Move Rope | 0.44 | 0.50 |
| Arrange Cabinet | 1.00 | 0.91 |
| Make Sandwich | 0.70 | 1.00 |
| Install Drywall | 0.50 | 1.00 |
Self-Reflection Analysis
As shown in Figure 1, the self-reflection component is one of the three key elements of our proposed metacognitive learning module. An example is shown in Figure 4 to demonstrate the workflow of metacognitive inference and self-reflection modules. To evaluate its practical impact, we introduce an additional metric in our experiments, the Reflection Success Rate, defined as the proportion of successful plan regenerations among all reflection attempts within task rounds that ultimately succeeded.
@@
As shown in the Table 4, in the Arrange Cabinet task, the framework achieves a perfect reflection success rate of , indicating that every initial failure was successfully recovered through self-reflection. In the Make Sandwich task, which involves long-horizon dependencies and stacking constraints, the framework recovers from of failed plans via self-reflection. Even in the most challenging Move Rope task, characterized by tight spatial coordination, the framework achieves a reflection success rate, underscoring the role of self-reflection in enabling meaningful plan recovery under physical constraints.
In our newly introduced Install Drywall task, the framework also demonstrates strong self-reflection capabilities. Specifically, using GPT-4 achieves an ideal reflection success rate of , signifying perfect recovery from all initial failures through reflection-driven adjustments. With LLaMA-3.1, while slightly lower, our framework still achieves a notable reflection success rate of , reinforcing the practical effectiveness of our metacognitive module in managing complex spatial coordination and real-time alignment corrections essential for this task.
These findings highlight the critical role of the self-reflection component within our metacognitive learning module. They also show that the proposed framework not only supports LLM-powered agents in initial plan generation but also enables them to proactively reason about and reflect on execution failures, providing the essential capability of reliable recovery and adaptation in zero-shot settings.
Discussions and Limitations
Our experimental results demonstrate that our metacognitive learning framework significantly enhances the effectiveness and adaptability of LLM-powered robots in zero-shot planning. Across all tasks, including the newly proposed Install Drywall benchmark, the integration of skill decomposition, metacognitive inference, and self-reflection leads to higher task success rates and fewer replans. Notably, in the Move Rope task, our framework exhibits creative behavior by adapting grasp points to handle IK failures and avoid collisions. These adaptations, while deviating from ground-truth plans, remain valid and effective, highlighting the role of metacognition in enabling structured creativity. Our statistical analysis also confirms that our framework not only improves overall success but also facilitates graceful recovery from failure. Finally, we observe that even an open-source LLM, such as LLaMA-3.1, performs competitively with GPT-4 under our framework, which suggests that the gains stem from the structure of our method rather than the scale or proprietary nature of the foundation model.
While our initial exploration demonstrates strong zero-shot performance and emergent structured creativity, the findings are mainly based on empirical observations. A formal framework is needed to analyze how metacognitive learning enables creativity, and future work will incorporate ablation studies to isolate component contributions, which is currently challenging due to the framework’s early-stage nature.
Conclusions